OpenStudy (anonymous):

PLEASE HELP!!!!!!! SO CONFUSED! Find the cube roots of 27(cos 279° + i sin 279°).

OpenStudy (kropot72):

Do you know De Moivre's Theorem?

OpenStudy (anonymous):

Yes (cosx +isinx) = cox(nx)) + isin(nx)

OpenStudy (kropot72):

Good! Then we can express the question as: \[(27\times (\cos 279+i \sin 279))^{\frac{1}{3}}=27^{\frac{1}{3}}\times (\cos 279+i \sin 279)^{\frac{1}{3}}\] Can you see what to do now?

OpenStudy (anonymous):

you would then divide 27 by 3? and 279 by 3? this is where i get very confused!

OpenStudy (kropot72):

You can't apply De Moivre's Theorem to the cube root of 27. \[27^{\frac{1}{3}}=\sqrt[3]{27}=?\]

OpenStudy (anonymous):

that would be 3 ?

OpenStudy (kropot72):

That's correct! So now we have: \[3\times (\cos 279+i \sin 279)^{\frac{1}{3}}\] and can apply De Moivre's Theorem.

OpenStudy (anonymous):

Okay, so where do i got from there?

OpenStudy (kropot72):

\[3\times (\cos 279+i \sin 279)^{\frac{1}{3}}=3(\cos \frac{279}{3}+i \sin \frac{279}{3})\]

OpenStudy (anonymous):

279/3 is 93, so thats my cubed root?

OpenStudy (kropot72):

The cube root is 3(cos 93 + i sin 93)

OpenStudy (anonymous):

ohhh okay, thats my final form! I need to do some more practicing! Thank you so much!!!

OpenStudy (kropot72):

You're Welcome :)